what is the probability of drawing 5 face cards from a 52 card deck? without replacement.
Answer by swincher4391(1107)

There are 12 face cards (J,Q,K)
On the first draw: 12/52
2nd: 11/51 and so on
5th: 8/48
We multiply the probabilities: (12/52)(11/51)(10/50)(9/49)(8/48) = 33/108290 = .0003